Three Piggies Enterprises has no debt. Its current total value is $76.8 million. Assume debt proceeds are used to repurchase equity.

Ignoring taxes, what will the company’s value be if it sells $35.4 million in debt? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)

Value of the firm $

Suppose now that the company’s tax rate is 35 percent. What will its overall value be if it sells $35.4 million in debt? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)

Value of the firm $
Answer #1

Value of levered firm = Value of unlevered firm + debt*tax rate

=76.8+35.4*0=76.8=76800000

Value of levered firm = Value of unlevered firm + debt*tax rate

=76.8+35.4*0.35=89.19=89190000